OpenStudy (anonymous):

solve the following systems of equations. if the system cannot be solved state that it cannot be solved and explain why? 3x+y=7 3x+y=10 i did 3x+y=7 --> y=-3x+7 3x+y=10 ---> y=_3x+10

OpenStudy (anonymous):

i have no idea what to do

OpenStudy (anonymous):

3x + y = 7 3x + y = 10 there is something to notice about these two equations right off the bat, if you solve for y as you did: y = -3x + 7 y = -3x + 10 you notice that both equations have a slope of -3, with a different y-intercept (7 and 10, respectively). if you graph both of these, you notice the lines will never touch. there won't ever be an answer for (x,y), cause there will never be an x and y were the lines intersect. the lines are parallel, and forever run past each other having the same slope

OpenStudy (anonymous):

you can not find any value of x and y that simultaneously satisfy both equations , well explained by funinbox

OpenStudy (anonymous):

thank u so much u explaineld it in a way where i actualli got it